OpenStudy (anonymous):

Suppose that 5,000 dollars is deposited in a savings account paying 6% annual interest compounded continuously. How much interest will be earned in eight years? the back of the book has the answers and its 3080.37 in interest.. but i dont know how to get there, please show work!!!! thanks!!!!

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

your the man help me out!!!!!!

OpenStudy (anonymous):

i think i have to use the equation S=Pe^rt maybe?

OpenStudy (gorv):

just use formula of cmpound interst

OpenStudy (anonymous):

what do i plug in though like what do i solve for

OpenStudy (gorv):

@mathhelp3242323

OpenStudy (anonymous):

got that equation but what does the P stand for? like what do i plug in

OpenStudy (anonymous):

and the n stand for?

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

interest compounded in a year 2 time s 3 times

OpenStudy (anonymous):

whatt lol

OpenStudy (gorv):

like interest is compounded quarterly

OpenStudy (anonymous):

can u wrtie it out.. like how to set it up please

OpenStudy (gorv):

sorry sorry sorry sorry ur formula is right

OpenStudy (gorv):

as interest compounded continuously so ur formula will be used for that

OpenStudy (anonymous):

no worries! lol

OpenStudy (gorv):

can solve further

OpenStudy (gorv):

@mathhelp3242323

OpenStudy (anonymous):

what do i put into S=Pe^rt though.. like thats where im stuck

OpenStudy (gorv):

judt plug in p r and t

OpenStudy (anonymous):

what do they stand for thoughh! lol like is p the 5000 r the .06 and t the 8?

OpenStudy (anonymous):

i got like 8080.37 but in the back of the book ists.. 3080.37

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

e =2.718

OpenStudy (gorv):

yeah lol that is total amount

OpenStudy (gorv):

for interest subtract the amount deposited 8080.37-5000=3080.37

OpenStudy (gorv):

hahahaha

OpenStudy (anonymous):

holy pellet hahahaa

OpenStudy (gorv):

hahahaah

OpenStudy (anonymous):

*pellet

OpenStudy (anonymous):

*pellet

OpenStudy (anonymous):

i keep typing pellet and it comes out as pellet... hahaha

OpenStudy (anonymous):

s h i t

OpenStudy (gorv):

gotta go now buddy

OpenStudy (anonymous):

Initally 200 milligrams of radioactie substance was present. After 6 hours, the mass had decreased by 3% . Construct an exponential model A(t)=A_0e^kt for the amount remainging of the decaying substance after t hours . Find the amount remaining after 24 hours?

OpenStudy (anonymous):

alright man thanks for all the help!